The comparative statements of financial position of Mikos Inc. as at December 31, 2017 and 2018, and its statement of earningMIKOS INC. Statement of Earnings For the Year Ended December 31, 2018 $ 883,000 Sales $441,000 5,100 34,100 232,000 13,100 CoRequired: 1. Prepare a statement of cash flows for Mikos Inc. for the year ended December 31, 2018 by using the indirect methNet cash flow used for investing activities 0 Cash flows from financing activities: Repayment of long-term note Net cash flowThe comparative statements of financial position of Mikos Inc. as at December 31, 2017 and 2018, and its statement of earnings for the year ended December 31, 2018, are presented below: MIKOS INC. Comparative Statements of Financial Position December 31 2018 2017 Assets Cash $ 10,100 $ 18,100 Short-term investments 63,700 34,300 Accounts receivable 73,500 30,400 Inventories, at cost 56,200 41,100 Prepaid expenses 5,100 8,200 Land 49,400 75,500 Property, plant, and equipment, net 284,400 184,500 Intangible assets 25,100 30,200 $ 567,500 $ 422,300 Liabilities and Shareholders’ Equity Accounts payable $ 18,100 $ 42,200 Income tax payable 8,200 2,100 Accrued liabilities 11,100 -0- Long-term notes payable 121,000 172,000 Contributed capital 222,000 65,500 Retained earnings 187,100 140,500 $ 567,500 $ 422,300 MIKOS INC. Statement of Earnings For the Year Ended December 31, 2018 Sales $ 883,000 Cost of sales $ 441,000 Amortization expense—intangible assets 5,100 Depreciation expense—property, plant, and equipment 34,100 Operating expenses 232,000 Interest expense 13,100 725,300 Earnings before income taxes 157,700 Income tax expense 47,310 Net earnings $ 110,390 Additional information is as follows: a. Land was sold for cash at its carrying amount. b. The short-term investments will mature in February 2019 (i.e. these are cash equivalents). c. Cash dividends were declared and paid in 2018. d. New equipment with a cost of $167,100 was purchased for cash, and old equipment was sold at its carrying amount. e. Long-term notes of $15,500 were paid in cash, and notes of $35,500 were converted to shares. Required: 1. Prepare a statement of cash flows for Mikos Inc. for the year ended December 31, 2018 by using the indirect method. When dealing with depreciation and amortization expenses (in the operating section), place them together as a single line item. (Negative answers should be indicated by a minus sign.)
Answer #1

Cash flow statement as follows:

MIKOS INC.
Statement of Cash Flows
For the Year Ended December 31, 2018
Cash flow from operating activities
Net Income $110,390
Adjustment for:
Add: Depreciation $34,100
Add: Amortization $5,100
Less: Increase in accounts receivables ($43,100)
Less: Increase in inventory ($15,100)
Add: Decrease in prepaid expenses $3,100
Less: Decrease in accounts payable ($24,100)
Add: Increase in income tax payable $6,120
Add: Increase in accrued liabilities $11,100 ($22,780)
Cash generated from operations $87,610
Cash flow from investing activities:
Sale of land $26,100
Purchase of equipment ($167,100)
Sale of equipment (184500+167100-284400) - Depr. 34100 $33,100
Net cash used in investing activities ($107,900)
Cash flow from Financing activities:
Payment of note ($15,500)
Common stock issued (222000-65500-35500) $121,000
Dividend paid (140500+110390-187100) ($63,790)
Net cash used in financing activities $41,710
Net Increase (decrease in cash) $21,420
Cash and short term investments beginning of the year $52,400
Cash and short term investments end of year $73,820

____________________________________________________________________________

Working as follows:

MIKOS INC.
Statement of Cash Flows
For the Year Ended December 31, 2018
Cash flow from operating activities
Net Income 110390
Adjustment for:
Add: Depreciation 34100
Add: Amortization 5100
Less: Increase in accounts receivables =30400-73500
Less: Increase in inventory =41100-56200
Add: Decrease in prepaid expenses =8200-5100
Less: Decrease in accounts payable =18100-42200
Add: Increase in income tax payable =8220-2100
Add: Increase in accrued liabilities 11100-0 =SUM(B72: B79)
Cash generated from operations =C70+C79
Cash flow from investing activities:
Sale of land =75500-49400
Purchase of equipment -167100
Sale of equipment (184500+167100-284400) - Depr. 34100 33100
Net cash used in investing activities =SUM(B83: B85)
Cash flow from Financing activities:
Payment of note -15500
Common stock issued (222000-65500-35500) 121000
Dividend paid (140500+110390-187100) -63790
Net cash used in financing activities =SUM(B89: B92)
Net Increase (decrease in cash) =C92+C86+C80
Cash and short term investments beginning of the year =18100+34300
Cash and short term investments end of year =SUM(C94: C95)